OpenStudy (anonymous):

How do i do this: X^2 + 2x + 3 = 0 Please show work!

OpenStudy (anonymous):

Which method do u want use?

OpenStudy (anonymous):

Quadratic

OpenStudy (anonymous):

\[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] for ax^2 +bx +c X^2 + 2x + 3 = 0 (compare) sub a, b, c in

OpenStudy (anonymous):

I already know the quadratic equation, theres something i am apparently not getting right in my work towards the end. Thats why i want to see the full working to the problem

OpenStudy (anonymous):

Show me what u have...

OpenStudy (anonymous):

U should have just asked someone to do it all for u in the question, it would have been much simpler....

OpenStudy (anonymous):

I misasked the question.. by " Please show work!" i meant to do it all

OpenStudy (anonymous):

What did you get when you plugged in a, b, and c?